a cube is sliding down an inclined plane inclined 'theta'to horizontal.suddenly it hits a small peg .find the angular velocity of the cube just after it hits the peg.its linear velocity is v.surface is frictionless. 

Grade:

1 Answers

Chetan Mandayam Nayakar
13 years ago

It can be shown(in case I am required to show, please ask me) that the moment of inertia about the axis of rotation is

(7/12)mR2, m is mass of cube and R is length of its side. Distance of centre of mass of cube from axis of

rotation =(R/√2)sin((π/4)+θ)  from conservation of angular momentum, we get (7/12)mR2w = (mvR/√2)sin((π/4)+θ)

w = angular velocity = √2(6v/7R)sin((π/4)+θ)